A ladder leaning against a house makes an angle of 30 degrees with the ground. the foot of the ladder is 7 feet from the foot of the house. How long is the ladder to the nearest foot?
Found 3 solutions by mananth, dabanfield, ikleyn:
Answer by mananth(16949)

A ladder leaning against a house makes an angle of 30 degrees with the ground. the foot of the ladder is 7 feet from the foot of the house. How long is the ladder to the nearest foot?
Cos 30 = base / hypotenuse
0.154= 7/hypotenuse
7/0.154
=45.45



Answer by dabanfield(803)

A ladder leaning against a house makes an angle of 30 degrees with the ground. the foot of the ladder is 7 feet from the foot of the house. How long is the ladder to the nearest foot?
Let L be the length of the ladder. The ladder and the side of the house form a right triangle with the ladder being the hypotenuse and the house wall being the side opposite the 30 degree angle. We have a theorem that says that in a right triangle the side opposite a 30 degree angle is half the hypotenuse, in this case L/2.
So by Pythagorean Theorem we have:
L^2 = (L/2)^2 + 7^2 or
L^2 = (L^2)/4 + 49
Solve the above for L.

Answer by ikleyn(53875)

.
A ladder leaning against a house makes an angle of 30 degrees with the ground.
the foot of the ladder is 7 feet from the foot of the house. How long is the ladder to the nearest foot?
~~~~~~~~~~~~~~~~~~~~~~~~~~~~


        The solution by @mananth is incorrect: it uses wrong value of cosine.
        I came to bring a correct solution.


Cos(30) = base / hypotenuse
0.866= 7/hypotenuse
hypotenuse = 7/0.866 = 8.08 meters, approximately.         ANSWER

Solved correctly.